OpenStudy (anonymous):

Calculus I Why is the antidervative of x^ n = [x^(n+1)]/ (n+1) + C]? i.e if I have x^2 the derivative is 2x why is the integral of 2x = [(x^2)/2] + C ? If I differentiate the (x^2)/2 I have to use the quotient rule and I don't get 2x which is what I started with. Quotient Rule (f'g-fg')/g^2

OpenStudy (anonymous):

I read the definition but it doesn't make sense to me. The antidervative is supposed to undo the derivative.

OpenStudy (anonymous):

\[ \int x\ dx=\frac{1}{2}x^2+C\\ \int 2x\ dx=2\int x\ dx = 2\cdot\frac{1}{2}x^2+C=x^2+C \]

OpenStudy (anonymous):

\[ \int ax^n\ dx=\frac{a}{n+1}x^{n+1}+C \] You're just dropping a 2 somewhere and that's why your results are getting mixed up.

OpenStudy (anonymous):

the dervative of \[x^2 = 2x\] so why do we divided by 2

OpenStudy (anonymous):

or multiply by 1/2

OpenStudy (anonymous):

nvm I found where I messed up. Thanks for all the help.